40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 16:46:31 2019, gautam, Update, IOO, IMC not locking PCdrive_RMS.png
    Reply  Tue Jun 4 10:56:31 2019, gautam, Update, IOO, IMC diagnostics IMCdiag.pdf
       Reply  Thu Jun 6 22:11:53 2019, Koji, Update, IOO, IMC diagnostics 
          Reply  Thu Jun 13 18:01:18 2019, aaron, Update, IOO, IMC diagnostics 
             Reply  Fri Jun 14 13:10:00 2019, aaron, Update, IOO, IMC diagnostics 
                Reply  Mon Jun 17 12:37:16 2019, aaron, Update, IOO, IMC diagnostics 
                   Reply  Tue Jun 18 20:35:07 2019, aaron, Update, IOO, IMC diagnostics 
                      Reply  Wed Jun 19 19:12:51 2019, aaron, Update, IOO, IMC diagnostics PMC_Error_Spectrum.pdfPDH_signal.pdfPDH_signal_full.pdf
                         Reply  Mon Jun 24 15:49:05 2019, aaron, Update, IOO, IMC diagnostics sweep_config_updated.png
                            Reply  Tue Jul 2 19:36:18 2019, aaron, Update, IOO, IMC diagnostics PDH_error.pdfPMC_Error_Spectrum.pdf
          Reply  Fri Jun 21 19:36:26 2019, Koji, Update, IOO, IMC diagnostics IMC_error.pdf
             Reply  Sun Jun 23 08:09:53 2019, gautam, Update, IOO, NPRO diagnostics consolidatedOscopeScreenCaps.pdfRINcomp.pdfPLL_OLTF.pdfPLLnoise.pdf
                Reply  Sun Jun 23 09:36:32 2019, gautam, Update, IOO, IMC is locking normally again PMCdiag.pdf
                   Reply  Sun Jun 23 14:43:14 2019, Koji, Update, IOO, IMC is locking normally again 
                   Reply  Mon Jun 24 08:12:10 2019, gautam, Update, IOO, IMC is locking normally again IMCdutycycle.png
                      Reply  Mon Jun 24 11:48:35 2019, gautam, Update, IOO, IMC in-loop error spectra and OLTF IMCfreqNoise.pdfIMC_OLTF.pdfIMC_CMboard.jpg
Message ID: 14647     Entry time: Mon Jun 3 16:46:31 2019     Reply to this: 14653
Author: gautam 
Type: Update 
Category: IOO 
Subject: IMC not locking 

Since ~ 2 hours ago, the IMC autolocker has not been able to keep the IMC locked. I don't see any obvious trends in the wall StripTool that may point to what's going on. For the brief periods in which a TEM00 mode is locked, the PC Drive RMS level is ~5x what the nominal level is, and while the autolocker is trying to lock the IMC, the PC drive RMS level is hovering around 4V DC, which is high. The PMC Error and Control signal spectra show huge 60 Hz (and harmonics) peaks, and indeed this is visible in the time domain signals as well (on ndscope or on the oscilloscope on the PSL table), but this is not a new feature in the last two hours. Usually, this kind of problem signals that either/both the c1psl or c1iool0 slow machines need to be power-cycled, but I confirmed that both machines are online and telnet-able. Possibilities: (i) some card in the c1psl / c1ioo crates have failed or (ii) something in the MC/FSS electronics chain has failed or (iii) there is a huge amount of excess high-frequency noise from the NPRO.

I am leaving the PSL shutter closed.

Attachment 1: PCdrive_RMS.png  29 kB  Uploaded Mon Jun 3 17:53:36 2019  | Hide | Hide all
PCdrive_RMS.png
ELOG V3.1.3-